Question 2 How much heat is required when 25 mols of 1-Hexene is heated at atmospheric pressure from 412 K to 1293.14 K?
Answer #1

heat capacity of hexane is 2.26 Kj/kg

mass of 25 mol of hexane = (25*86) =2150 g = 2.15kg where 86 is the molar mass of hexane

heat required = m*Cp*(1293.14 - 412)

= 2.15*2.26*881.14

= 4281.5 KJ